Simplify this expression. (b+4)^2 divided by b^2+7b+12 I don't know what to do with the bottom part. I believe the top would be (b+4)(b+4) but not sure of the rest.
Answer by jorel1380(3719)

(b+4)^2 divided by b^2+7b+12=
(b+4)(b+4)/(b+4)(b+3)=b+4/b+3..